OpenStudy (s):

Find lim as x-> pi ((pi-x)^2)/(sin^2x) ?

OpenStudy (anonymous):

\[ \lim_{x\to \pi}\frac{(\pi-x)^2}{\sin^2(x)} \]Is this correct?

OpenStudy (s):

correct

OpenStudy (anonymous):

Do you know l'Hospital's rule?

OpenStudy (s):

as far as I remember, I should take derivative of the top and bottom, right?

OpenStudy (anonymous):

yes

OpenStudy (s):

so I will get 2(pi-x) / cos^2x ?

OpenStudy (anonymous):

sorry, I was distracted...

OpenStudy (s):

it's OK

OpenStudy (anonymous):

\[ d((\pi-x)^2) = 2(\pi-x)d(\pi-x) = -2(\pi-x)dx \]So the top would be: \[ 2x-2\pi \]

OpenStudy (anonymous):

\[ d(\sin^2(x)) = 2\sin(x)d(\sin(x)) = 2\sin(x)\cos(x)dx =\sin(2x)dx \]So the bottom is: \[ \sin(2x) \]

OpenStudy (anonymous):

This still leads to the indeterminate form of \(0/0\), so we have to do l'Hospital a second time.

OpenStudy (anonymous):

Does it make sense how I got the derivatives?

OpenStudy (s):

yes, I just checked online, thank you

OpenStudy (s):

I took derivative one more time and got -1/cos 2x which is -1

OpenStudy (anonymous):

....

OpenStudy (anonymous):

\[ d(2x-2\pi) = 2dx \]Top derivative is \(2\).

OpenStudy (anonymous):

\[ d(\sin(2x)) = \cos(2x)d(2x) = 2\cos(2x)dx \]Bottom derivative is \(2\cos(2x)\).

OpenStudy (s):

right.. I mistakenly put 2pi- 2x

OpenStudy (anonymous):

So final answer?

OpenStudy (s):

1

OpenStudy (s):

thank you very much! Wish you all the best!